# New AF Einstein–Maxwell instantons

### v1

@@ Abstract
We explicitly construct a three–parameter family of asymptotically flat Einstein–Maxwell instantons. These solutions are toric, regular, and free of conical and orbifold singularities on the manifold $M=\mathbb{CP}^{2}\setminus S^{1}$. They are counterexamples to the Euclidean Einstein–Maxwell Black Hole Uniqueness Conjecture. In the case of vanishing charge these instantons reduce to the Chen–Teo Ricci flat instantons.
@@ I Introduction
It has remained an open problem, noted in [8] and [23], to construct an Einstein-Maxwell generalisation of the Chen-Teo gravitational instanton, which would at the same time address the possible violation of the Euclidean version of the charged Lorentzian black hole uniqueness theorems. This issue was also explicitly formulated as an open problem by Yau [26, Problem 116]. In this note we settle this question, and construct a three-parameter family of smooth, regular, AF Einstein-Maxwell instantons, which reduce to the Chen–Teo instantons on the boundary of the parameter space. We first construct a six-parameter family of asymptotically locally flat (ALF) Einstein-Maxwell structures: the corresponding metrics are asymptotic to circle bundles over $S^{2}$, which fail to be trivial if the asymptotic NUT charge does not vanish. These solutions have conical singularities and generalise the vacuum metrics found in [8]. We then demonstrate that the parameters can be adjusted to obtain a new AF Einstein-Maxwell gravitational instanton, i.e. a smooth solution free of orbifold and conical singularities and with zero asymptotic NUT charge on the manifold $M=\mathbb{CP}^{2}\setminus{S^{1}}$. We find a three-parameter family of such spaces, and we furthermore show that they are conformally Kähler. It then follows that the Weyl tensor is one-sided type $D$, which implies that the solutions do not admit a Lorentzian section—thus there is no contradiction with Lorentzian uniqueness theorems.
@@ III AF Einstein–Maxwell instantons
The asymptotic NUT charge is proportional to the $\tau$–component of ${l}_{1}-{l}_{4}$. Thus, the AF condition requires ${l}_{1}={l}_{4}$ (since the $\phi$-component of ${l}_{1}-{l}_{4}$ vanishes). From the above regularity requirements, we have ${l}_{1}=-\epsilon{l}_{3}+pl_{2}$ (with $p=\epsilon p^{\prime}$) and ${l}_{4}=-\epsilon^{\prime}{l}_{2}+q{l}_{3}$. So ${l}_{1}-{l}_{4}=0$ gives $p=-\epsilon^{\prime}$ and $q=-\epsilon$. Therefore, the combination of smoothness and asymptotic flatness is equivalent to
@@ IV Asymptotic charges and limits
and defining $\tau=c\,T+{l}_{1}^{\tau}\Phi$, $\phi={l}_{1}^{\phi}\Phi$, $x=r_{2}-\frac{r_{2}c\sqrt{k}}{R}\cos^{2}(\frac{\Theta}{2})$, $y=r_{2}+\frac{r_{2}c\sqrt{k}}{R}\sin^{2}(\frac{\Theta}{2})$. Then when $R\to\infty$, the metric (1) satisfies
$g={\rm d}T^{2}+{\rm d}R^{2}+R^{2}({\rm d}\Theta^{2}+\sin^{2}{\Theta}\,{\rm d}\Phi^{2})+O(R^{-1}).$ (9)
(The leading term in (9) with $(T,\Phi)$ $2\pi$–periodic is the flat metric on $S^{1}\times\mathbb{R}^{3}$. The new coordinate system is found by identifying the two semi–infinite rod directions $l_{1}=l_{4}$ with $\partial/\partial\Phi$). We can extract the mass $m$ and angular momentum $j$ from the sub-leading terms in (9), i.e. from the asymptotic expansions $g_{TT}=1-\frac{2m}{R}+O(R^{-2})$ and $g_{T\Phi}=\frac{2j\sin^{2}\Theta}{R}+O(R^{-2})$ [18]: we find
@@
The Maxwell field (2) has the asymptotic behavior
@@
The solution (1) with $Q=0$ is the Ricci–flat five-parameter family constructed by Chen and Teo [8]. The corresponding Maxwell field (2) is self–dual, so it does not contribute to the energy–momentum tensor. For general $Q$ the asymptotic norm of $\partial_{\tau}$ is finite as long as $1<\nu<1-2Q$. The boundary values $\nu=-1$ and $\nu=1-2Q$ correspond to ALE (rather than ALF or AF) metrics. If $\nu=-1$ the ASD part of the Maxwell field vanishes and the metric $g$ belongs to the Gibbons–Hawking class with three centres in the harmonic potential lying on one axis. This is a result of a calculation where we find that $\lim_{\nu\rightarrow-1}(H_{0}/H)=(1-Q)^{-1}$. We similarly expect that $\nu=1-2Q$ is the charged Plebański–Demiański metric.

### v2

# New asymptotically flat Einstein–Maxwell instantons

@@ Abstract
We disprove the Euclidean Einstein–Maxwell Black Hole Uniqueness Conjecture, and thus demonstrate that the semi-classical properties of coupled gravitational and electromagnetic fields are more subtle than expected from Lorentzian general relativity, where the Kerr—Newman family of metrics yields the most general stationary and asymptotically flat black holes with a single event horizon. This is achieved by an explicit construction of a new three–parameter family of asymptotically flat Einstein–Maxwell instantons. These solutions are toric, regular, and free of conical and orbifold singularities on the manifold $M=\mathbb{CP}^{2}\setminus S^{1}$. In the case of vanishing charge, these instantons reduce to the Chen–Teo Ricci flat instantons.
@@ I Introduction
It has remained an open problem, noted in [8] and [23], to construct an Einstein-Maxwell generalisation of the Chen-Teo gravitational instanton, which would at the same time address the possible violation of the Euclidean version of the charged Lorentzian black hole uniqueness theorems. This issue was also explicitly formulated as an open problem by Yau [26, Problem 116]. In this note we settle this question, and construct a three-parameter family of smooth, regular, AF Einstein-Maxwell instantons, which reduce to the Chen–Teo instantons on the boundary of the parameter space. We first construct a six-parameter family of asymptotically locally flat (ALF) Einstein-Maxwell structures: the corresponding metrics are asymptotic to circle bundles over $S^{2}$, which fail to be trivial if the asymptotic NUT charge does not vanish. These solutions have conical singularities and generalise the vacuum metrics found in [8]. We then demonstrate that the parameters can be adjusted to obtain a new AF Einstein-Maxwell gravitational instanton, i.e. a smooth solution free of orbifold and conical singularities and with zero asymptotic NUT charge on the manifold $M=\mathbb{CP}^{2}\setminus{S^{1}}$. We find a three-parameter family of such spaces and, in addition, show that they are conformally Kähler. It then follows that the Weyl tensor is one-sided type $D$, which implies that the solutions do not admit a Lorentzian section—thus there is no contradiction with Lorentzian uniqueness theorems.
@@ III AF Einstein–Maxwell instantons
The asymptotic NUT charge which we compute in the next section is proportional to the $\tau$–component of ${l}_{1}-{l}_{4}$. Thus, the AF condition requires ${l}_{1}={l}_{4}$ (since the $\phi$-component of ${l}_{1}-{l}_{4}$ vanishes). From the above regularity requirements, we have ${l}_{1}=-\epsilon{l}_{3}+pl_{2}$ (with $p=\epsilon p^{\prime}$) and ${l}_{4}=-\epsilon^{\prime}{l}_{2}+q{l}_{3}$. So ${l}_{1}-{l}_{4}=0$ gives $p=-\epsilon^{\prime}$ and $q=-\epsilon$. Therefore, the combination of smoothness and asymptotic flatness is equivalent to
@@ IV Asymptotic charges and limits
and defining
$\tau=$ $c\,T+\frac{1}{2}({l}_{1}^{\tau}+{l}_{4}^{\tau})\Phi,\quad\phi=\frac{1}{2}({l}_{1}^{\phi}+{l}_{4}^{\phi})\Phi$
$x=$ $r_{2}-\frac{r_{2}c\sqrt{k}}{R}\cos^{2}(\frac{\Theta}{2})\quad y=r_{2}+\frac{r_{2}c\sqrt{k}}{R}\sin^{2}(\frac{\Theta}{2})$
so that $\frac{1}{2}(l_{1}+l_{4})=\partial/\partial\Phi$. Then, when $R\to\infty$, the metric (1) satisfies
$g=$ $({\rm d}T+2n\cos{\Theta}{\rm d}\Phi)^{2}+{\rm d}R^{2}$
$+R^{2}({\rm d}\Theta^{2}+\sin^{2}{\Theta}\,{\rm d}\Phi^{2})+O(R^{-1}).$ (9)
The NUT parameter $n$ can be computed from the asymptotic form of the metric (1) as
$n=\frac{\sqrt{k}}{2c(r_{1}-r_{2})(r_{2}-r_{3})(r_{3}-r_{4})}(n_{0}+n_{1}Q),$ (10)
where
$n_{0}=$ $(1+\nu^{2})(r_{2}^{3}+r_{1}r_{3}r_{4})$
$+2\nu r_{2}(r_{1}r_{2}+r_{2}r_{3}+r_{2}r_{4}-r_{1}r_{3}-r_{1}r_{4}-r_{3}r_{4})$
$n_{1}=$ $(1-\nu)(r_{1}-r_{2})(r_{2}-r_{3})(r_{2}-r_{4}).$
In the Ricci–flat case $Q=0$ our expression for $n$ agrees with that of Chen–Teo [8]. A non–zero NUT charge obstructs the asymptotic flatness, and the existence of globally defined Euclidean time $T$ coordinate such that the orbits of $\partial_{T}$ form a trivial $S^{1}$ bundle at infinity.
If condition (6) holds then $l_{4}=l_{1}$ and $n=0$. The leading term in (IV) with $(T,\Phi)$ $2\pi$–periodic is the flat metric on $S^{1}\times\mathbb{R}^{3}$. We can now extract the mass $m$ and angular momentum $j$ from the sub-leading terms in (IV), i.e. from the asymptotic expansions $g_{TT}=1-\frac{2m}{R}+O(R^{-2})$ and $g_{T\Phi}=\frac{2j\sin^{2}\Theta}{R}+O(R^{-2})$ [18]: we find
@@
The Maxwell field (2) has an asymptotic behavior
@@
The solution (1) with $Q=0$ is the Ricci–flat five-parameter family constructed by Chen and Teo [8]. The corresponding Maxwell field (2) is self–dual, so it does not contribute to the energy–momentum tensor. For general $Q$ the asymptotic norm of $\partial_{\tau}$ is finite as long as $1<\nu<1-2Q$. The boundary values $\nu=-1$ and $\nu=1-2Q$ correspond to ALE (rather than ALF or AF) metrics. If $\nu=-1$ the ASD part of the Maxwell field vanishes and the metric $g$ belongs to the Gibbons–Hawking class with three centres in the harmonic potential lying on one axis. This is a result of a calculation where we find that $\lim_{\nu\rightarrow-1}(H_{0}/H)=(1-Q)^{-1}$. The other ALE limit $\nu=1-2Q$ appears to be more subtle. If $Q=0$, then it reduces to the Plebański–Demiański (PD) metric [8]. We expect that $\nu=1-2Q$ gives the charged PD metric, but we have not been able to verify it directly, nor to argue that the hidden symmetry transformation we used to find (1) reduces in the limit to a transformation relating the charged and uncharged PD metrics.